OpenStudy (anonymous):

integrate (e^x)/(square root(e^x+1)) dx

OpenStudy (anonymous):

use the u-substitution .

OpenStudy (anonymous):

make u=e^x+1 du=e^x dx

OpenStudy (anonymous):

so you have integral of sqaure root of u. and hten integrate that. 2/3 (e^x+1)^3/2

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):